# and Reggio Emilia, Italy

### v2

@@ 3. Bento
For a comprehensive evaluation of the dataframe libraries on general, real-world data preparation use cases, we selected four datasets covering a variety of sizes, complexities, and features: (i) Athlete  [^20] https://www.kaggle.com/datasets/heesoo37/120-years-of-olympic-history-athletes-and-results, collecting 0.2M records about the results achieved by the athletes through 120 years of Olympics; (ii) Loan  [^21] https://www.kaggle.com/datasets/wordsforthewise/lending-club, containing 2M rows about loan applicants and their financial profiles from the LendingClub company; (iii) Patrol  [^22] https://www.kaggle.com/datasets/faressayah/stanford-open-policing-project, composed of 27M records about 11 years of traffic stops by California law enforcement agencies (Pierson et al., 2020); (iv) Taxi  [^23] https://www.kaggle.com/competitions/nyc-taxi-trip-duration, containing 77M records about taxi trips in New York City in 2015. A summary of their main features is displayed in Table 3.
Collected from Kaggle, all of these datasets had previously been adopted in related work about data preparation and dataframe libraries. The three largest datasets were indeed used to evaluate the performance of Modin (Petersohn et al., 2021), due to the variety of their features. Instead, Athlete was exploited to analyze how different preparators are supported by commercial data preparation tools (Hameed and Naumann, 2020).
@@
For each dataset, we selected from Kaggle the most voted notebook that proposed a non-trivial solution for the task at hand (i.e., Kaggle competitions) and extracted its data preparation pipeline—typically the first part of the notebook, preceding the definition of a machine learning model.
@@
  • Exploratory data analysis (EDA): preparators for supporting the exploratory analysis of data features, to better understand the data at hand and detect errors or anomalies.
@@
Unless differently stated, we load data from CSV files and measure the execution time under three key settings: (i) function-core, which evaluates each preparator alone (if the library adopts lazy evaluation, it requires to force the execution after each API call); (ii) pipeline-stage, which measures the runtime for the execution of each stage (i.e., I/O, EDA, DT, and DC); (iii) pipeline-full, which measures the runtime for the execution of the entire pipeline. Finally, our framework also supports the pipeline execution in Docker containers, which allow to specify core and memory usage and isolate Python environments to prevent package conflicts.
@@
All measurements are obtained as the average value over ten runs on a machine equipped with two AMD EPYC Rome 7402 CPUs (48 threads running at 2.8 GHz) and 512 GB of RAM, using Python 3.9, when not specified otherwise. The machine also features a NVIDIA A100 GPU with 40 GB of RAM and CUDA 12.1. The Ray engine was configured with default settings, resulting in 48 worker threads running. The Dask engine was configured comparably, leading to the creation of 8 workers and 48 threads for each execution. To ensure accuracy and avoid warm-up overhead, the assessment of execution time occurs once the system (e.g., JVM) has completed its warm-up process. To assess the performance of the libraries across machines with different hardware specifications, we simulated three distinct machine configurations, as detailed in Table 4. Finally, we evaluated scalability on incremental samples of Taxi and Patrol.
@@ 4. Evaluation Results
In this section, we report and analyze the results of the extensive experimental evaluation of the presented dataframe libraries. Our main goal is to provide data scientists and practitiones with useful insights for supporting them in the selection of the best solution for their data preparation tasks. Thus, our comparison is designed to assess the performance of dataframe libraries based on the operations to carry out (considering both the distinct preparators and the benefits introduced by lazy evaluation, when supported), the size and the features of the dataset at hand, and the configuration of the machine on which the pipeline is executed.
@@
Figure 1. Average runtime for each stage of the data preparation pipeline (EDA, DT, DC) on different datasets.
@@ 4.1. Evaluation on Data Preparation Pipelines
Summary—For EDA, Polars is generally the best performer, up to 100 times faster than Pandas. For DT, if a GPU is available, CuDF generally outperforms other libraries. For DC, ModinR achieves notable results on the largest datasets. Finally, despite not supporting the Parquet format, DataTable is the fastest library for I/O.
For each of the selected datasets (Athlete, Loan, Patrol, Taxi), Figure 1 shows the runtime for the execution of three main stages of the data preparation pipeline: EDA, DT, and DC. The I/O stage is considered separately in Figures 4 and 4. Figure 1 considers each stage in its entirety, allowing to perform lazy evaluation at stage level when supported. Additionally, in Figure 2 we present the independent performance of each distinct preparator (i.e., we force the execution separately for each preparator), normalized with respect to Pandas (the de facto standard library for managing dataframes). For each preparator, the red line represents therefore Pandas as the baseline, while distinct markers denote the speedup for each library, which is defined as follows:
$speedup=\frac{Time{\langle\textrm{Pandas},\textit{{prep}}\rangle}}{Time{\langle\textit{{lib}},\textit{{prep}}\rangle}}$
where $Time{\langle\textit{{lib}},\textit{{prep}}\rangle}$ is the time required by the library lib to run the preparator prep. Thus, a value above the red line denotes that the library outperforms Pandas; otherwise, it takes more time than Pandas to run the preparator. The logarithmic scale enables to easily assess the order of magnitude of the improvement (if any) of each library over Pandas.
In the next subsections, we analyze in detail the results obtained for each of the main data preparation stages, considering both the single preparators and the entire stages. Note that not all preparators have the same impact on the pipeline, but their weight depends on the time required for their execution (as we point out when analyzing the related results).
@@
Figure 2. Speedup over Pandas for each preparator on different datasets.
Explorative Data Analysis (EDA)
For EDA, Polars clearly stands out as the best performer on all datasets but Patrol, where CuDF outperforms other dataframe libraries (Figures 1(a)-1(d)).
In this stage, four preparators alone account for 95% of the total execution time: locate missing values (isna), locate outliers (outlier), search by pattern (srchptn), and sort values (sort). Note that Polars generally registers the best performance for two of them, locate missing values and locate outliers (Figure 2), where it is up to ${10^{4}}$ times faster than Pandas and 0.6 times faster than the second best performer, DataTable. In fact, Polars can quickly locate missing values by exploiting the specific metadata stored by Arrow arrays [^24] https://pola-rs.github.io/polars/user-guide/expressions/null/. Nevertheless, DataTable is the fastest in locating missing values on Patrol (Figure 2(c)), with a speedup of ${10^{4}}$ over Pandas and ${10^{3}}$ over Polars, thanks to its implementation of binary-search-based filtering. As highlighted by Table 3, Patrol is mostly composed of string attributes. For this type of data, Polars uses Arrow large_utf8 buffers, which are more expensive to filter than Python strings/chars, requiring therefore extra work.
For the other two most relevant preparators, in Figure 2(c) we can observe that Vaex excels in searching by pattern, being about ${10^{3}}$ times faster than Pandas and 10 times faster than the second best performer, CuDF. Vaex tends to be very efficient in column-wise operations due to its internal columnar organization of the data. On the other hand, it is much less efficient when applying a row-wise preparator, for instance in locating outliers, where it is even about 100 times slower than Pandas. Conversely, when locating missing values, Vaex shows relatively good performance on dense datasets such as Taxi and Patrol, though it is among the less efficient on sparse datasets like Loan, where its performance is does not provide improvements over Pandas (Figure 2(b)).
For sorting values, CuDF and Polars are the fastest libraries overall. SparkSQL shows remarkable advantages over Pandas as the datasets grow larger, while SparkPD maintains similar performance to the baseline. Note that in Athlete, Modin converts by default the dataframe back and forth to Pandas for the sorting operation on multiple columns, hence in this case it is the worst performer, up to 100 times slower than Pandas.
Finally, for the remaining preparators, get dataframe statistics (stats) and query columns (query), Modin outperforms other libraries on larger datasets exploiting its multithreading power, which even makes ModinD 10 times faster than Polars when running the entire stage on Taxi. Overall, Pandas maintains good performance throughout the stage.
@@
Figure 3. Average runtime for reading CSV and Parquet files.
@@
Figure 4. Average runtime for writing CSV and Parquet files.
Data Trasformation (DT)
In this stage, CuDF emerges as the best performer on Loan (Figure 1(f)) and Patrol (Figure 1(g)), while SparkSQL and Polars register the best performance (also thanks to lazy evaluation) on Athlete (Figure 1(e)) and Taxi (Figure 1(h)), respectively. The most time intensive preparators in this stage are categorical encoding (catenc), calculate new column (apply), and group dataframe (group). For catenc and group, CuDF always shows the best performance, being at least up to 100 times faster than the second best performer (Polars and ModinR, respectively). Once again, Vaex consistently outperforms the other libraries on column-wise operations (e.g., apply), being ${10^{3}}$ times faster than the second best performer CuDF (Figure 2(d)). On the other hand, this causes a drastic drop in its performance when moving to operations such as group, join, or pivot.
As in the previous stage, SparkPD results as the worst performer on several preparators. This is mainly due to the latency introduced by translating Pandas operations into Spark execution plans. Since SparkSQL does not require this translation, it often achieves significantly better performance.
When datasets scale up, libraries like ModinD and ModinR, designed for resource-intensive applications, exhibit superior performance due to their optimized handling of large-scale data compared to their performance on smaller datasets.
@@
When moving to data cleaning operations, CuDF consistently outperforms other libraries on all datasets, except for Taxi (Figure 1(l)), where Vaex stands out as the best performer (with remarkable results achieved by Polars and Modin with both engines).
The pipelines designed for the smallest datasets (Athlete and Loan) include a preparator, deduplicate rows (dedup), accounting for 70% of the overall execution time for the stage. Pandas register good results on these datasets, especially on Athlete, where only CuDF performs better (Figure 1(i)). On Loan, it is possible to notice how Vaex equals or (even significantly) exceeds the performance of CuDF for multiple preparators (Figure 2(b)), but the greater overall impact of the ones where it is outperformed by CuDF (especially the great difference in deduplicating rows) leads to a consistent difference in the overall execution time for this stage (Figure 1(j)).
Results change when moving to larger datasets (Figures 1(k) and 1(l)). CuDF excels on Patrol, where it is over ${10^{3}}$ times faster than Pandas for both considered preparators, i.e., delete empty rows (dropna) and change date & time format (chdate), slightly outperforming ModinR. Nevertheless, it struggles on Taxi, where Vaex achieves the best results instead, followed closely by Modin (with both engines) and Polars.
@@
Figure 5. Average runtime for the execution of the entire data preparation pipelines.
@@
Loading data represents a fundamental step for the success of any data-driven process. In particular, the CSV format has established itself as the de facto standard for reading and writing raw data, and Pandas is recognized as an efficient and effective tool for its ingestion (Vitagliano et al., 2023). Nowadays, the diffusion of Arrow is pushing towards an increasing adoption of the Parquet format [^25] https://parquet.apache.org, known for its efficient data compression. In particular, developers often load Parquet data into memory and convert it to Arrow (Liu et al., 2023), recognized as an ideal in-memory transport layer for such data [^26] https://arrow.apache.org/docs/python/parquet.html.
@@
Figures 4 and 4 show performance in read and write operations on both file formats, CSV and Parquet. DataTable, which maps the file into memory and navigates through it using pointers, records the best overall performance in reading CSV files, but it does not support the Parquet format. Polars is the best performer in reading Parquet files due to its capability to load them directly into memory using Arrow. When dealing with small datasets like Athlete, Parquet and CSV often exhibit almost the same performance, unlike Polars and SparkSQL (Figure 3(a)). In some cases (e.g., Vaex or CuDF), the use of Parquet can even lead to worse performance, due to the need for decompressing the data. When the size increases, the use of Parquet generally produces better results; for instance, reading Parquet files with Polars is over 100 times faster than reading CSV files (Figure 3(d)).
Write operations exhibit a wide range of performance, mainly due to the need of compressing data before saving. Parquet files come with additional metadata and encoding, which may introduce some overhead for the writing process. Nevertheless, Parquet performs better than CSV across all datasets as it integrates various efficient compression and encoding approaches, while CSV is plain text and uncompressed (Liu et al., 2023).

### v3

@@ 3. Bento
For a comprehensive evaluation of the dataframe libraries on general, real-world data preparation use cases, we selected four datasets covering a variety of sizes, complexities, and features: (i) Athlete (Kaggle, 2018), collecting 0.2M records about the results achieved by the athletes through 120 years of Olympics; (ii) Loan (Kaggle, 2019a), containing 2M rows about loan applicants and their financial profiles from the LendingClub company; (iii) Patrol (Kaggle, 2019b), composed of 27M records about 11 years of traffic stops by California law enforcement agencies (Pierson et al., 2020); (iv) Taxi (Kaggle, 2017), containing 77M records about taxi trips in New York City in 2015. A summary of their main features is displayed in Table 2. Note that we compute the percentage of missing values as the number of null cells over all cells in the dataframe. Collected from Kaggle, all of these datasets had previously been adopted in related work about data preparation and dataframe libraries. The three largest datasets were indeed used to evaluate the performance of Modin (Petersohn et al., 2021), due to the variety of their features. Instead, Athlete was exploited to analyze how different preparators are supported by commercial data preparation tools (Hameed and Naumann, 2020).
@@
Preparator SparkPD SparkSQL Modin Polars CuDF Vaex DataTable
I/O load dataframe (read) ✓✓ ✓✓ ✓✓ ✓✓
output dataframe (write) ✓✓ ✓✓ ✓✓ ✓✓
EDA locate missing values (isna) ✓✓ $\circ$ ✓✓ ✓✓ $\circ$
locate outliers (outlier) ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
search by pattern (srchptn) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
sort values (sort) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
get columns list (getcols) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
get columns types (dtypes) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
get dataframe statistics (stats) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
query columns (query) ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
DT cast columns types (cast) ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
delete columns (drop) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
rename columns (rename) ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
pivot table (pivot) ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
calculate column using expressions (calccol) ✓✓ $\circ$ ✓✓ ✓✓ $\circ$ ✓✓ $\circ$
join dataframes (join) ✓✓ $\circ$ ✓✓ ✓✓ $\circ$ $\circ$
one hot encoding (onehot) ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ $\circ$
categorical encoding (catenc) ✓✓ ✓✓ ✓✓ $\circ$
group dataframe (group) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
DC change date & time format (chdate) ✓✓ ✓✓ $\circ$ ✓✓ $\circ$ $\circ$
delete empty and invalid rows (dropna) ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
set content case (setcase) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
normalize numeric values (norm) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
deduplicate rows (dedup) ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
fill empty cells (fillna) ✓✓ ✓✓ $\circ$ ✓✓ ✓✓ $\circ$
replace values occurrences (replace) ✓✓ ✓✓ $\circ$ ✓✓ $\circ$
edit & replace cell data (edit) ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ ✓✓
For each dataset, we selected three notebooks among the top-voted Kaggle entries that proposed a non-trivial solution for the task at hand (i.e., Kaggle competitions) and extracted its data preparation pipeline—typically the first part of the notebook, which precedes the definition of a machine learning model.
@@
  • Exploratory data analysis (EDA): preparators to support the exploratory analysis of data features, to better understand the data at hand and detect errors or anomalies.
@@
Unless differently stated, we load data from CSV files and measure the execution time under three key settings: (i) function-core, which evaluates each preparator alone (if the library adopts lazy evaluation, it requires to force the execution after each API call); (ii) pipeline-stage, which measures the runtime for the execution of each of the four stages (i.e., I/O, EDA, DT, and DC); (iii) pipeline-full, which measures the runtime for the execution of the entire pipeline. Finally, our framework also supports the execution of the pipeline in Docker containers, which allow to specify core and memory usage and isolate Python environments to prevent package conflicts.
@@
All measurements are obtained as the average value over ten runs on a machine equipped with two AMD EPYC Rome 7402 CPUs (48 threads running at 2.8 GHz) and 512 GB of RAM, using Python 3.9, when not specified otherwise. The machine also features a NVIDIA A100 GPU with 40 GB of RAM and CUDA 12.1. The Ray engine was configured with default settings, resulting in 48 worker threads running. The Dask engine was configured comparably, leading to the creation of 8 workers and 48 threads for each execution.
To ensure accuracy and avoid warm-up overhead, the assessment of execution time occurs once the system (e.g., JVM) has completed its warm-up process. To assess the performance of the libraries across machines with different hardware specifications, we simulated three distinct machine configurations, as detailed in Table 4. Finally, we evaluated scalability on incremental samples of Taxi and Patrol.
@@ 4. Evaluation Results
In this section, we report and analyze the results of the extensive experimental evaluation of the presented dataframe libraries. Our main goal is to provide data scientists and practitiones with useful insights for supporting them in the selection of the best solution for their data preparation tasks. Therefore, our comparison is designed to assess the performance of dataframe libraries based on the operations to carry out (considering both the distinct preparators and the benefits introduced by lazy evaluation, when supported), the size and the features of the dataset at hand, and the configuration of the machine on which the pipeline is executed.
@@
Figure 1. Average speedup over Pandas computed for each stage (EDA, DT, DC) on the three data preparation pipelines.
@@ 4.1. Evaluation on Data Preparation Pipelines
Summary—For EDA, Polars is generally the best performer. For DT, if a GPU is available, CuDF generally outperforms other libraries. For DC, Vaex achieves notable results on the largest datasets. Finally, CuDF and Polars appear to be the best choices to read and write CSV files, respectively. For each of the selected datasets (Athlete, Loan, Patrol, Taxi), Figure 1 and Figure 2 show the average speedup over Pandas achieved by its alternatives in the execution of the three data preparation pipelines per dataset, focusing on the stages of EDA, DT, and DC. In particular, Figure 1 considers each stage in its entirety, allowing to perform lazy evaluation at the stage level when supported, Figure 2 shows the performance separately for each distinct preparator (i.e., we force the execution for each of them). The I/O stage is considered separately in Figures 4 and 4. The bars in Figure 1 and the markers in Figure 2 denote the speedup achieved over Pandas (the red line), defined as follows:
$\textit{{speedup}}=\frac{Time{\langle\textit{{Pandas}},\textit{{prep/stage}}\rangle}}{Time{\langle\textit{{lib}},\textit{{prep/stage}}}\rangle}$
[…]
$\textit{{Impact}}=\frac{Time{\langle\textit{{dataset}},\textit{{prep}}\rangle}}{Time{\langle\textit{{dataset}},\textit{{stage}}\rangle}}\times 100$
where $Time{\langle\textit{{dataset}},\textit{{prep}}\rangle}$ is the average time for each stage preparator and $Time{\langle\textit{{dataset}},\textit{{stage}}\rangle}$ is the sum of the average times for all preparators of the stage on the three dataset pipelines [^5] To minimize bias from particularly fast or slow libraries, both in the numerator and in the denominator time values below the 20th percentile and above the 80th percentile were excluded..
In the next subsections, we analyze in detail the results obtained for each of the main data preparation stages, considering both the single preparators and the entire stages.
@@
Figure 2. Average speedup over Pandas computed for each preparator of the three data preparation pipelines per datasets. For each preparator, we report the number of calls in each pipeline and its average impact on the stage execution time.
@@
Figure 3. Average runtime for reading CSV and Parquet files (lower is better).
@@
Figure 4. Average runtime for writing CSV and Parquet files (lower is better).
Explorative Data Analysis (EDA)
For EDA, Polars clearly stands out as the best performer on all datasets (Figures 1(a)-1(d)). As depicted in Figure 2, Polars always registers the best performance for isna and outlier, the preparators with the greatest impact on Athlete and Loan, respectively. For the former preparator, Polars is up to 10,000 times faster than Pandas and 3 times faster than the second-best performer, i.e., DataTable. Both libraries avoid element-wise comparisons by using a special encoding to efficiently track null values: while Polars relies on the validity bitmap used by most types of Arrow arrays in their metadata (Polars, 2024c), DataTable encodes them with sentinel values (H2O, 2024). For the latter preparator, outliers are located by filtering the dataframe using the values returned by the quantile function. Pandas relies on NumPy exact quantile computation (involving sorting and interpolation) (Pandas, 2024c), while Polars and Spark employ approximate methods that avoid sorting (Apache Spark, 2024b). The performance gap between SparkSQL and SparkPD arises from differences in their underlying implementations.
Although Vaex tends to be very efficient in column-wise operations (e.g, srchptn) and can handle filtering well by internally tracking selected rows without copying data, operations based on percentiles are penalized by the complexity of the required calculations (i.e., determining min/max column values, cumulative sums, and grid interpolation) (Vaex, 2024b).
For sort, CuDF and Polars are the fastest libraries overall. While CuDF leverages on high-performance C++ parallel algorithms from the Thrust library (NVIDIA, 2024b), Polars achieves high performance through its efficient multi-threaded Rust implementation (Polars, 2024e). SparkSQL shows remarkable advantages over Pandas as the datasets grow larger, while SparkPD performs similarly to the baseline. Modin performs significantly worse than Pandas on Athlete as it requires to apply Pandas implementation of the sorting preparator within each data partition (Modin, 2024d), introducing significant latency to partition and merge results. For stats, DataTable outperforms Pandas by up to 50 times by computing statistics either during the creation of the Frame object or efficiently on-demand when needed, while Modin (with both Ray and Dask) is up to 200 times faster than the baseline on Taxi thanks to its multi-threading capabilities.
Vaex achieves notable performance for query, for which it excels on the Taxi dataset and is only outperformed by Polars, which enables fast querying by leveraging mask operations (Polars, 2024d), on the smaller ones. Finally, Pandas generally maintains overall good performance for getcols and dtypes, where PySpark is much slower due to the overhead associated to its inherently distributed architecture, optimized for complex tasks rather than simple ones.
Data Transformation (DT)
In this stage, CuDF and Polars emerge as the best performers overall, highlighting the benefits of GPU acceleration and Rust code optimizations, respectively. In particular, CuDF outperforms Polars on Loan (Figure 1(f)) and Patrol (Figure 1(g)), while the opposite is true on Athlete (Figure 1(e)). As depicted in Figure 2, CuDF always achieves the best results for catenc, while for onehot and group it is the best performer on all datasets but Athlete, where Polars outperforms the alternatives. SparkSQL shows excellent results on Patrol (Figure 1(g)): despite not excelling for single preparators (Figure 2(c)), lazy evaluation provides remarkable benefits. Moreover, it clearly outperforms SparkPD, which results as the worst performer on several preparators, mainly due to the latency introduced by translating Pandas operations into Spark execution plans.
Vaex optimizations lead to outstanding performance on column-wise preparators (e.g., calccol and drop), but also to drastic drops moving to operations such as group, join, or pivot. Modin, designed for resource-intensive applications, improves its performance as datasets scale up, with ModinR being the best performer for pivot on Taxi. DataTable registers remarkable performance on that preparator too, despite not supporting it natively (DataTable, 2024a), but particularly excels for cast, due to direct memory manipulation and in-place casting (DataTable, 2024b). On the other hand, libraries relying on the Arrow columnar format (e.g., Polars), with its safety check and abstraction layers, may experience additional overhead (Polars, 2024a).
@@
When it comes to DC, Polars achieves the best results on Athlete (Figure 1(i)), outperforming CuDF (which performs better for all preparators if considered separately, as depicted in Figure 2(a)) thanks to lazy evaluation. CuDF excels on Loan (Figure 1(j)) thanks to its performance with dedup, which has by far the highest runtime among the preparators, thus deserving an in-depth analysis in the following. CuDF implements dedup natively, by using factorization to identify duplicates (Pandas, 2024b). Differently, Polars tracks and manages the presence or absence of unique values using bitmasking (Polars, 2024d), while SparkSQL uses grouping and aggregations to find duplicate rows (Apache Spark, 2024c). Finally, Vaex and DataTable lack an official implementation for dedup.
Results change when moving to larger datasets (Figures 1(k) and 1(l)), with Vaex achieving the best results both on Patrol (followed closely by CuDF and ModinR) and on Taxi (with remarkable performance by Polars and both Modin versions). On Patrol, Vaex is the best performer on both preparators (Figure 2(c)): dropna, where it exploits its efficient filtering mechanisms, and chdate, for which it relies on NumPy operations (faster compared to more complex operations found in other libraries) to enhance conversion and manipulation of columns (Vaex, 2024c).
@@
Figure 5. Average speedup over Pandas for the execution of the entire data preparation pipelines.
@@
Loading data represents a fundamental step for the success of any data-driven process. In particular, the CSV format is considered as the de facto standard for reading and writing raw data, and Pandas is recognized as an efficient and effective tool for its ingestion (Vitagliano et al., 2023). Nowadays, the diffusion of Arrow is pushing towards an increasing adoption of the Parquet format (Apache Parquet, 2024), known for its efficient data compression. In particular, developers often load Parquet data into memory and convert it to Arrow (Liu et al., 2023), recognized as an ideal in-memory transport layer for such data (Apache Arrow, 2024b).
@@
Figures 4 and 4 show performance in reading and writing both file formats, CSV and Parquet. When reading small datasets like Athlete, Parquet and CSV often exhibit similar performance, with exceptions such as Polars or SparkSQL (Figure 3(a)). In some cases (e.g., Vaex or CuDF), the use of Parquet can even lead to worse performance, due to the need for decompressing the data. When the size increases, the use of Parquet generally produces better results; for instance, reading Parquet files with Polars is over 100 times faster than reading CSV files (Figure 3(d)).
Overall, CuDF and Vaex appear to be the best performers in reading CSV files. Vaex is particularly fast thanks to chunked reading, conversion to optimized formats (e.g., HDF5), and efficient use of memory (Vaex, 2024a). CuDF, relying on GPU acceleration and efficient memory management, maintains similar performance for both file formats. DataTable, which maps the file into memory and navigates through it using pointers, performs good in reading CSV files, but it does not support Parquet. Polars and Vaex excel in reading Parquet files thanks to their capability to load them directly into memory using Arrow.
Write operations exhibit a wide range of performance. Overall, Polars and CuDF are the top performers and Parquet performs better than CSV across all datasets as it integrates efficient compression and encoding approaches, while CSV is plain text and uncompressed (Liu et al., 2023). Nevertheless, the overhead introduced by compression and encoding operations can penalize Parquet in some scenarios, as highlighted by CuDF on small datasets (e.g., Athlete and Loan). Further, Polars presents a reported issue in writing Parquet files, which determines slower performance and larger file size (Polars, 2024d). Finally, Modin (with both engines) demonstrates the best performance on Parquet, particularly for large datasets, by leveraging parallel processing across multiple cores (Modin, 2024a).
4.2. Impact of Lazy Evaluation
{mdframed}
SummaryCuDF generally outperforms other libraries when running the entire pipeline, while SparkSQL and ModinR achieve remarkable results on all datasets. Lazy evaluation often brings relevant benefits, improving performance by 20% on average (up to almost 80%) over its eager counterpart. Figure 5 reports the average speedup over Pandas achieved by its alternatives for the execution of the entire data preparation pipelines on each dataset. For libraries supporting lazy evaluation (i.e., SparkPD, SparkSQL, and Polars), we also point out the difference compared to eager evaluation.
Considering the overall performance on the entire pipeline, CuDF generally outperforms other libraries by effectively exploiting the power of GPU, with the only exception of Athlete (Figure 5(a)), where Polars is the best performer. PySpark (except for SparkPD on Athlete), ModinR, and Polars show very solid performance overall, as well as Vaex on the largest datasets. Pandas performs relatively well on the smallest datasets, but it is outperformed by all alternatives on larger ones, due to its eager approach causing high memory consumption (Sinthong and Carey, 2021).
Polars generally achieves good results with both kinds of evaluations, as its eager API is highly optimized and in many cases it internally relies on its lazy counterpart before immediately collecting the result. Lazy evaluation leverages techniques such as streaming processing, early filtering, and projection pushdown, achieving a performance improvement of up to 25% on Patrol (Figure 5(c)). On one hand, it can be noticed how long logical plans can limit the optimizations introduced by lazy evaluation. Yet, the execution plan overhead of SparkSQL does not yield substantial improvements for either smaller or larger datasets, with an improvement of 40% on Taxi and almost no changes in performance on Loan. Libraries like SparkPD, leverage to Pandas-specific optimizations (e.g., vectorized operations and efficient memory management) that translate for instance into a 80% performance improvement on Patrol (Figure 5(c)).
4.3. Scalability
{mdframed}
SummarySparkSQL is the best library for scaling to large datasets on a single machine. Indeed, for both Patrol and Taxi, it is the only one able to execute the entire pipelines with the laptop configuration. Polars follows closely, but requires a lot of resources in terms of RAM, exhibiting the worst scalability on machines with limited resources. In this section, we assess how dataframe libraries scale by varying the size of the dataset and the configuration of the underlying machine. In particular, we take into account three different machine configurations, whose CPU and RAM are incrementally increased as described in Table 4 (which also reports the specifications of the Dask and Ray engines used by Modin) to reproduce typical configurations for laptops, workstations, and servers.